Today's work involved the continued excavation of the unexcavted portion of AB1 (locus 12), the final removal of the AB3 locus 1 topsoil matrix and stone cleaning, the commencement of the AB2 locus 10 excavations, and the continued mapping of the hill with Ian's robotic total station.In AB1 we continued to excavate the locus 12 cultural fill containing moderate amounts of materials and significant charcoal flecking. We eventually reached the continuation of the locus 13 cobble wall fall locus, and there are several in situ ceramic and bone objects visible. We also uncovered the continuation of the locus 3 wall running west into operation AB2. Tomorrow we will attempt to reach the wall fall locus across the entirety of locus 12. In AB2 we began excavation of locus 10 sterile fill and hit what may be bedrock in some spots. Material densities are very low, so it's not surprising that we may have reached a bedrock layer. We need to define what the relationship is between that locus and the wall fall, dense materials matri directly south (locus 9). In AB3 we finished cleaning the locus 1 rock fall in trench north. Tomorrow we will draw and photograph locus one before proceeding with the excavation of the pale yellow clayey silt of locus 2. Mapping proceeded on the first and second north terraces. Benchmark 1 appears to have visibilty from the east side of the hill to the northwest side. We will definitely need another benchmark on the west side in order to complete the map.